OpenStudy (anonymous):

Find (fgh)(-2) if f(x) = x2 - 5, g(x) = 3x + 1, h(x) = x + 3. fgh(-2)=

OpenStudy (sheraz12345):

firstly Substitute -2 in H(x) and then substitute the answer in g(x) and then that answer in f(x)

OpenStudy (anonymous):

h(x) =1

OpenStudy (sheraz12345):

yes put that in g(x)

OpenStudy (anonymous):

f(x)= -1 right ?

OpenStudy (sheraz12345):

Yes @alyygirl

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

g(x) = -5

OpenStudy (sheraz12345):

That is it -1 is ur answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im also confused about this

OpenStudy (sheraz12345):

What is the confusion?

OpenStudy (anonymous):

hold on

OpenStudy (sheraz12345):

ok

OpenStudy (anonymous):

do you understand that ?

OpenStudy (sheraz12345):

I'm trying to .. Juz wait

OpenStudy (anonymous):

ok thanks

zepdrix (zepdrix):

Ooo interesting problem =o

OpenStudy (anonymous):

yea Its weird huh

zepdrix (zepdrix):

\[\Large\rm g(\color{red}{h(-1)})\]Your inner function is being evaluated at x=-1. So trace the red line over to x=-1, what is the output value? the y-value?

OpenStudy (sheraz12345):

h(-1 ) = -3 Ii guess

OpenStudy (sheraz12345):

and g(-3) = 3

OpenStudy (sheraz12345):

So answer will be 3 for first part

OpenStudy (anonymous):

just 3 for (gh)(-1)=3?

OpenStudy (sheraz12345):

We know y=f(x) So we have are given with x which is -1 .. From graph we find value of y for h(-1) and so on

OpenStudy (sheraz12345):

Yes @alyygirl

OpenStudy (anonymous):

ok and (f h g) (-4) =

OpenStudy (anonymous):

-12 times f?

OpenStudy (sheraz12345):

How -12 ?

OpenStudy (anonymous):

-3 times 4?

OpenStudy (anonymous):

whats f?

OpenStudy (sheraz12345):

at h(-4) we have the value y=2 then g(2)=3 and f(3)=-4 so i guess -4 is the answer .. i hope i am not wrong

OpenStudy (anonymous):

-4 for the second part? and 3 the first part

OpenStudy (anonymous):

ill trust your answer better then mine lol

OpenStudy (sheraz12345):

lOL Yea... -4 for 2nd part and 3 for first... Is this some sort of test or quiz

OpenStudy (anonymous):

its online math class

OpenStudy (anonymous):

no practice

OpenStudy (anonymous):

it sucks

OpenStudy (sheraz12345):

Ahhh okay Good luck then .. and lemme know if my answers were correct

OpenStudy (anonymous):

ok I will :)

OpenStudy (anonymous):

you there?

zepdrix (zepdrix):

He is offline :c sup?

OpenStudy (anonymous):

i dont understand this

zepdrix (zepdrix):

Understand what? The graph problem with the composition stuff?

zepdrix (zepdrix):

lolol did you take a picture of your computer screen? :3 it's all tilted. dat's funny.

OpenStudy (anonymous):

yea cuz you cant type it haha

OpenStudy (anonymous):

I wish this was easy omg

zepdrix (zepdrix):

So yah, piece-wise functions can be a little weird. It's one of the three pieces, depending on where our x is at that given moment. So when our x is -2, which piece do we want to use? (Look at the intervals on the right)

OpenStudy (anonymous):

the second one

zepdrix (zepdrix):

So the first and second piece both mention -2. But it looks like the first piece has the `or equal to`, doesn't it?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Ok so when x=-2, our function becomes:\[\Large\rm f(\color{orangered}{x})=(\color{orangered}{x})^2+4\]

zepdrix (zepdrix):

\[\Large\rm f(\color{orangered}{-2})=(\color{orangered}{-2})^2+4\]

OpenStudy (anonymous):

f(-2)=8

zepdrix (zepdrix):

yay!

OpenStudy (anonymous):

thats it? or no

zepdrix (zepdrix):

Yes that solves your first blank.

OpenStudy (anonymous):

nice not bad

zepdrix (zepdrix):

Any confusion on how our function simply became: \(\Large\rm f(\color{orangered}{x})=(\color{orangered}{x})^2+4\) ?

zepdrix (zepdrix):

When x is AT or below -2, we use that specific piece to evaluate.

zepdrix (zepdrix):

What do you think for the next one? :U

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

f (3) 2x+3?

zepdrix (zepdrix):

Check your intervals carefully. You seem to keep wanting to go to that middle one for some reason.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

If x=3, that means x is larger than 1, right? It's not between -2 and 1.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

shoot ok so f(3)= 2x(3)+3

OpenStudy (anonymous):

Im guessing this is wrong?

zepdrix (zepdrix):

Yes :( You went to the middle piece again........

zepdrix (zepdrix):

The BOTTOM PIECE corresponds to x values larger than 1, yes?

OpenStudy (anonymous):

ok yea

OpenStudy (anonymous):

so its square root x+1 equation we use

zepdrix (zepdrix):

Yes.

OpenStudy (anonymous):

f(3) = square root3x +1

zepdrix (zepdrix):

I guess you're having a little trouble with function notation. Try to understand the colors I'm using here, So when \(\Large\rm x\gt1\),\[\Large\rm f(\color{orangered}{x})=\sqrt{\color{orangered}{x}+1}\]\[\Large\rm f(\color{orangered}{3})=\sqrt{\color{orangered}{3}+1}\]See how we're REPLACING each orange with 3? The x's get `replaced` with 3. You shouldn't have a 3x anywhere.

OpenStudy (anonymous):

ok 12?

OpenStudy (anonymous):

3+1 =4*3=12

zepdrix (zepdrix):

where is the 4*3 step coming from?

OpenStudy (anonymous):

adding inside the the square root and multiplying 3?

zepdrix (zepdrix):

Did you read my comment that I made? You should not having a 3x anywhere. You should only have a 3. `THE 3 IS REPLACING YOUR X`

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

it 2

zepdrix (zepdrix):

yay good job!

OpenStudy (anonymous):

ok good :)

OpenStudy (anonymous):

so thats the second box

OpenStudy (anonymous):

now for the last one we use the middle?

OpenStudy (anonymous):

equation

zepdrix (zepdrix):

Yes we FINALLY get to use the middle one ;) lol

zepdrix (zepdrix):

The interval for the middle piece INCLUDES -1, see the `or equal to` sign on the 1?

OpenStudy (anonymous):

lol

zepdrix (zepdrix):

So yes middle.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Not -1, I mean 1, sorry.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

is the answer for the last 0